Please do not delete this question.
Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents ½g, the gravitational pull due to gravity (measured in feet per second 2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.


How high will the ball be above ground after 1 second?

Found 2 solutions by stanbon, bucky:
Answer by stanbon(75887)

Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents ½g, the gravitational pull due to gravity (measured in feet per second^2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
---------------------
s(t) = -16t^2+32t+0
------------------------
How high will the ball be above ground after 1 second?
s(1) = -16(1^2) + 32(1)
s(1) = -16 + 32
s(1) = 16 ft.
The height will be 16 ft after 1 second.
====================
Cheers
Stan H.

Answer by bucky(2189)

Since the problem states that the baseball is initially shot upward from ground level, you
know that s0 = 0.
.
In that case the equation is reduced to:
.
s = -16t2 + v0t
.
The problem also tells you that the initial velocity is +32 ft/sec (the + means that it is
in the upward direction). Substitute this and the equation becomes:
.
s = -16t^2 + 32t
.
You are asked how high the ball is after 1 second. The units that apply in this equation are
feet and seconds. Substitute 1 for t and this will determine s in feet.
.
s = -16(1)^2 + 32(1) = -16 + 32 = 16 ft
.
The ball will be 16 feet high when the time after launch is 1 second.
.
Hope this helps you to understand the problem.